The curve y=2x^3-3x^2-36x+15 has a turning point at (3, -66) and (-2,59). Identify whether each of these points is a maximum or a minimum. (show all your work).


1
Expert's answer
2021-05-24T15:48:14-0400

Given curve-


y=2x^3-3x^2-36x+15


Differentiating y w.r.t. x-


"y'=6x^2-6x-36~~~~~-(1)"


Again differentiating it w.r.t x-


"y''=12x-6"


Putting x=3, "y''=12(3)-6=30>0"


Since y''>0, Point (3,-66) is a minimum points.



Now Putting x=-2, "y''=12(-2)-6=-30<0"


As y''<0, So Point (-2,59) is local maximum.